The Chartered Financial Analyst (CFA®) designation is fast becoming a requirement for serious investment professionals. Although it requires a successful completion of three levels of grueling exams, it also entails promising careers with lucrative salaries. A student of finance is curious about the average salary of a CFA®charterholder. He takes a random sample of 25 recent charterholders and computes a mean salary of $136,000 with a standard deviation of $35,000. Use this sample information to determine the 99% confidence interval for the average salary of a CFA® charterholder. Assume that salaries are normally distributed. Use Table 2. (Round intermediate calculations to 4 decimal places. Round "t" value to 3 decimal places and final answers to the nearest whole number.)
Answer #1

Ans:

sample size,n=25

df=25-1=24

critical t value=tinv(0.01,24)=2.797

99% confidence interval for average salary

=136000+/-2.797*(35000/sqrt(25))

=136000+/-19579

=(116421, 155579)